# Using dust to constrain dark matter models

### v1

@@ Abstract
In this paper, we use hydrodynamic zoom-in simulations of Milky Way-type haloes to explore using dust as an observational tracer to discriminate between cold and warm dark matter universes. Comparing a cold and 3.5keV warm dark matter particle model, we tune the efficiency of galaxy formation in our simulations using a variable supernova rate to create Milky Way systems with similar satellite galaxy populations while keeping all other simulation parameters the same. Cold dark matter, having more substructure, requires a higher supernova efficiency than warm dark matter to achieve the same satellite galaxy number. These different supernova efficiencies create different dust distributions around their host galaxies, which we generate by post-processing the simulation output with the powderday codebase. Analysing the resulting dust in each simulation, we find 4.5 times more dust in our cold dark matter Milky Way halos compared with warm dark matter. The distribution of dust out to R is then explored, revealing that the warm dark matter simulations are noticeably less concentrated than their cold dark matter counterparts, although differences in substructure complicate the comparison. Our results indicate that dust is a possible unique probe to test theories of dark matter.
@@
dark matter – galaxies: evolution – galaxies: formation – dust, extinction – software: simulations1 Introduction
@@
The leading description of the cosmology of the Universe, known as cold dark matter (CDM), combines a model for dark energy, , with cold dark matter, some non-interacting massive particles assumed to be non-relativistic during a significant fraction of the evolution of the universe. This description has been remarkably successful in explaining various phenomena in the large-scale Universe, including the angular power spectrum of the cosmic microwave background (Aghanim et al., 2020) and baryonic acoustic oscillations (Eisenstein et al., 2005). Nevertheless, despite numerous observations supporting the existence of dark matter (Zwicky, 1933; Rubin et al., 1985; Clowe et al., 2006), a dark matter particle has yet to be found, leaving its true nature as one of the great mysteries in science.
@@
There has been a vast array of experiments designed to detect an imprint of dark matter particles, both through direct and indirect detection methods. Altogether, these experiments have probed an enormous mass range, from eV up to TeV scales (see Klasen et al. (2015) for a review), though with limited sensitivity. So far, all these experiments have set stringent limits on the dark matter parameter space (mass and interaction rate with ordinary matter or a hypothetical dark sector), except for the DAMA experiment. They claimed the detection of an annual modulation signal compatible with a cold dark matter particle in the late 1990s (Bernabei et al., 2000), and maintain this claim as the DAMA/LIBRA experiment (Bernabei et al., 2008, 2021). This signal is in strong tension with the recent findings from the XenonnT (Aprile et al., 2023), LZ (Aalbers et al., 2023) and the plethora of other direct detection experiments, including those using a similar technique (Coarasa et al., 2024; Adhikari et al., 2021), which is difficult to explain theoretically. There is also a possible hint in favour of dark matter particles from the excess of gamma-rays from the galactic centre found in the FERMI-LAT data (Goodenough & Hooper, 2009; Hooper & Goodenough, 2011), which is also highly debated (see, e.g. Song et al. (2024) for the latest analysis). If confirmed, this would correspond to a 30GeV candidate with very suppressed interactions with nuclei to evade the direct detection limits. A dark matter discovery will take more than two highly debated hints (especially given the plethora of negative searches), so the best way to progress is to exploit the wealth of information coming from astrophysical observations. These observations are particularly useful to probe whether dark matter is collisionless (e.g. Boehm & Schaeffer, 2005; Boehm et al., 2014; Mosbech et al., 2023), and if so, determine the lower bound of the particle mass (e.g. Nadler et al., 2021a). The abundance of negative direct detection searches combined with the flexibility of fits to astronomical data allows for the possibility of a more exotic dark matter particle.
@@
The physical properties of different candidate dark matter particles manifest as macroscopic differences on astronomical scales, which are often expressed phenomenologically as cold and warm models of dark matter. To combat the difficulties in observing dark matter in the Universe, cosmological simulations have been used to quantify which characteristics of a particular dark matter model best fit the observations. Results from early cold dark matter simulations were able to shed light on several dark matter properties, such as the discovery of a universal halo density profile across a wide range of mass scales, and the abundance of intermediate to high mass haloes (Navarro et al., 1996; Springel et al., 2005). However, these early simulations also resulted in several well-known tensions with observations, for example, the missing satellites problem (Klypin et al., 1999; Moore et al., 1999), the core-cusp problem (de Blok, 2010), and the too-big-to-fail problem (Boylan-Kolchin et al., 2012), see Sales et al. (2022) for a recent review. Much research has been dedicated to resolving these tensions by improving the quality of the simulations (e.g. Crain et al., 2015; Kaviraj et al., 2017; Pillepich et al., 2018; Hernández-Aguayo et al., 2022; Schaye et al., 2023). These more advanced simulations in recent years bring higher resolutions with better force accuracy and, importantly, the inclusion and advancement of baryonic models.
@@
The missing satellites problem has been the subject of much examination. State-of-the-art CDM simulations of the 1990s and early 2000s overproduced low-mass dark matter haloes by an order of magnitude compared to expectations based on Milky Way satellite observations of the time. Alternative dark matter models were seen as one possible solution, in particular, warm dark matter. Warm dark matter models introduce a free-steaming scale below which dark matter density fluctuations are erased, thus reducing the number of haloes at high redshift and eventually the number of satellites around Milky Way-like galaxies (Bode et al., 2001; Avila-Reese et al., 2001; Boehm & Schaeffer, 2005). The scale at which these low-mass objects are suppressed is set by the mass of the dark matter particle. A heavier warm dark matter particle would look more similar to cold dark matter, whereas a lighter particle produces a much more significant suppression of low-mass halo formation.
@@
The other accepted solution to the missing satellites problem generally invokes baryonic physics to suppress star formation in low mass haloes, often using the feedback of supernova (Governato et al., 2007; Sawala et al., 2016). The energy generated by a supernova event is transferred back into the gas in the simulations, heating the gas. In low-mass haloes with shallow gravitational potential wells, supernova feedback can cause gas outflows with enough energy to eject gas out of the halo entirely, eventually halting star formation. In larger haloes, supernovae still drive outflows, but the potential wells are generally deep enough to retain the gas, which can be accreted back into the galaxy. More efficient supernova models111Throughout this text, we refer to ‘supernova efficiency’ as the stellar mass fraction converted to supernova, i.e. the parameter in Equation 1 of Springel & Hernquist (2003)., where stars capable of going supernova are formed at higher rates, are able to suppress star formation in more massive galaxies, though as galaxies get larger AGN become the dominant form of feedback (Croton et al., 2006). Alternatively, reionisation was proposed as a solution to the missing satellites problem (Bullock et al., 2000; Benson et al., 2002), where the photons heat the gas to such a point that low mass haloes can’t cool the gas to form galaxies. Importantly, suppression due to reionisation still allows the formation of low-mass galaxies, but preferentially for haloes that form prior to reionisation; thus, the reionisation simulation parameters can be tuned to match satellite observations.
@@
The tensions between simulations and observations have also been eased in recent years by more complete surveys of the satellite galaxies around the Milky Way (e.g. McConnachie, 2012; McConnachie & Venn, 2020). These surveys have been able to observe more low luminosity galaxies and galaxy remnants (streams). Combining the updated observations with increasingly precise modelling has led authors such as Sales et al. (2022) to conclude that the missing satellites problem can be fully explained within a cold dark matter universe.
@@
However, baryonic solutions to the missing satellites problem come with their own complications. The feedback mechanisms used to reduce low-mass galaxy formation are often conceptually similar but numerically inconsistent between simulations. Supernova feedback alone has an array of implementations across popular codebases, including simply increasing nearby gas temperature (Springel & Hernquist, 2003; Governato et al., 2007), to more complicated kinetic and thermal feedback descriptions (Dalla Vecchia & Schaye, 2008; Chaikin et al., 2023). In reality, there are many forms of feedback, from AGN and supernova to photo-heating, which all come with their own implementations within different simulation codes (for a recent review see Vogelsberger et al., 2020). While simulations using a CDM cosmology can solve the missing satellites problem through high efficiency feedback, the solutions are degenerate with each other and not mutually exclusive of warm dark matter. There is still work to be done to find methods to distinguish between dark matter models.
@@
In this paper, we explore a new method for distinguishing the underlying dark matter species in simulations. It is well known that the missing satellites problem can be solved by implementing either a warm dark matter with reduced feedback efficiency or cold dark matter with increased feedback efficiency. Therefore, we combine each dark matter type with different supernova efficiencies to create simulations with similar Milky Way satellite populations. However, supernovae are also critically important in the creation and transportation of dust (Aguirre, 1999; Todini & Ferrara, 2001; Bianchi & Schneider, 2007), and so these different supernova efficiencies are expected to produce different dust properties around the host halo. We explore if these different dust properties can then be used to infer (or at least discriminate) the underlying dark matter model. The techniques we present to analyse the dust provide a proof of concept that can be leveraged in future work with more in-depth analysis of dust in simulations and observations.
@@
This paper is organised as follows. In Section 2, we discuss the codes used and simulation setup. Section 3 explains the simulation results with a discussion following in Section 4. We conclude in Section 5. Throughout this paper we assume a flat Planck18 cosmology (Aghanim et al., 2020), where , , , , =0.840, and .
@@ 2.1 Simulation code setup
For the primary analysis, we perform zoom-in hydrodynamic simulations focusing on Milky Way-like systems. We run cold and warm dark matter, the latter having a mass of 3.5 keV. This is in tension with more stringent observational constraints on the warm dark matter mass (Viel et al., 2013; Nadler et al., 2021a, b; Dekker et al., 2022), but within the limits set by Benito et al. (2020) and Newton et al. (2021). The lighter mass is used to ensure an identifiable difference in the small-scale structure that is wholly attributed to the choice of dark matter.
@@
To generate initial conditions for each cold and warm dark matter simulation, we use music, the Multi-Scale Initial Conditions generator (Hahn & Abel, 2011). We apply the Eisenstein & Hu (1998) fitting function to create the input power spectra for the cold dark matter simulations and assume a truncated power spectrum for the warm dark matter simulations, using the Viel et al. (2005) transfer of the fitting function. The Viel et al. (2005) parameterisation of the transfer function gives us a warm dark matter half mode mass, M⊙, at which point we expect significant divergence between cold and warm dark matter simulations. This work is focused on relative differences, so the fitting functions are accurate enough compared to traditional Boltzmann solvers. music disregards any extra thermal velocity the warm dark matter particles may carry in the initial conditions, which can produce extra spurious haloes if not treated correctly (Klypin et al., 1993; Colín et al., 2008). Using this generalised approach to generate the warm dark matter power spectrum avoids any assumptions about the nature of the particle itself. The initial conditions for all simulations are created with the same random seed.
@@
We run the parent dark matter only and the subsequent zoom-in simulations with gadget-4 (Springel et al., 2021), an N-body and SPH code with significant improvements to older versions of gadget. gadget-4 makes use of a TreePM algorithm (Xu, 1995; Bagla, 2002) for the collisionless particles and a pressure-entropy SPH prescription (Springel & Hernquist, 2002) for the gas particles. The gravitational potential is softened with a cubic spline kernel and becomes Newtonian at distances , where epsilon is the simulation force resolution, and the SPH properties are smoothed over 64 nearest neighbours. The baryonic prescriptions implemented with the public release of gadget-4 are discussed in Section 2.3. Finally, we use a Friends-of-Friends algorithm (Davis et al., 1985) and subfind-hbt (Han et al., 2018; Springel et al., 2021) for halo and subhalo identification.
@@
Both parent and zoom-in simulations are run from redshift to , with a box side length of 65Mpc, noting that the zoom region is 0.1% the total volume. The parent simulations have 5123 equal mass dark matter particles, which equates to an approximate mass resolution of and a force resolution of 10kpc. The zoom-in simulations have a dark matter particle resolution of 2.4 M⊙ and a gas particle resolution of . More details about the simulations can be found in Table 1. The parent simulations are used to identify the haloes resembling the Milky Way halo to resimulate.
@@
| Box side length | 65Mpc | 65Mpc |
| Dark matter mass resolution | 5.68 M⊙ | 2.44 M⊙ |
| Baryon mass resolution | NA | 3.99 M⊙ |
| Stellar mass resolution | NA | 3.99 M⊙ |
| Force resolution | 10kpc | 500pc |
@@ 2.2 Target selection
The halo catalogues produced by subfind-hbt (Springel et al., 2021) are used to identify haloes similar in mass to the Milky Way in the parent cold and warm dark matter simulations. We select haloes with a mass M between 1-2 M⊙, where M is the mass enclosed with a radius R such that the density is 200 times the critical density of the universe, . This is consistent with observed Milky Way masses from Bland-Hawthorn & Gerhard (2016). We then apply an isolation criterion to these haloes. We exclude any potential resimulation target if there is another halo with a mass greater than 0.5 M⊙ within 8Mpc. This ensures that the Milky Way analogue is the largest object in the resimulated region, and that we aren’t investigating a galaxy undergoing a major merger. We use the same criteria and check both warm and cold dark matter parent simulations to verify the target halo exists in both simulations with similar properties. If there are multiple haloes that fit these criteria, we use the halo closest to the centre of the box.

### v2

@@ Abstract
In this paper, we use hydrodynamic zoom-in simulations of Milky Way-type haloes to explore using dust as an observational tracer to discriminate between cold and warm dark matter (WDM) universes. Comparing a cold and 3.5 keV WDM particle model, we tune the efficiency of galaxy formation in our simulations using a variable supernova rate to create Milky Way systems with similar satellite galaxy populations while keeping all other simulation parameters the same. Cold dark matter (CDM), having more substructure, requires a higher supernova efficiency than WDM to achieve the same satellite galaxy number. These different supernova efficiencies create different dust distributions around their host galaxies, which we generate by post-processing the simulation output with the powderday codebase. Analysing the resulting dust in each simulation, we find 4.5 times more dust in our CDM Milky Way haloes compared with WDM. The distribution of dust out to is then explored, revealing that the WDM simulations are noticeably less concentrated than their CDM counterparts, although differences in substructure complicate the comparison. Our results indicate that dust is a possible unique probe to test theories of dark matter.
@@
software: simulations – dust, extinction – galaxies: evolution – galaxies: formation – dark matter1 Introduction
@@
The leading description of the cosmology of the Universe, known as cold dark matter (CDM), combines a model for dark energy, ,, with CDM, some non-interacting massive particles assumed to be non-relativistic during a significant fraction of the evolution of the universe. This description has been remarkably successful in explaining various phenomena in the large-scale Universe, including the angular power spectrum of the cosmic microwave background (Aghanim et al., 2020) and baryonic acoustic oscillations (Eisenstein et al., 2005). Nevertheless, despite numerous observations supporting the existence of dark matter (Zwicky, 1933; Rubin et al., 1985; Clowe et al., 2006), a dark matter particle has yet to be found, leaving its true nature as one of the great mysteries in science.
@@
There has been a vast array of experiments designed to detect an imprint of dark matter particles, both through direct and indirect detection methods. Altogether, these experiments have probed an enormous mass range, from eV up to TeV scales (see Klasen et al. (2015) for a review), though with limited sensitivity. So far, all these experiments have set stringent limits on the dark matter parameter space (mass and interaction rate with ordinary matter or a hypothetical dark sector), except for the DAMA experiment. They claimed the detection of an annual modulation signal compatible with a CDM particle in the late 1990s (Bernabei et al., 2000), and maintain this claim as the DAMA/LIBRA experiment (Bernabei et al., 2008, 2021). This signal is in strong tension with the recent findings from the XenonnT (Aprile et al., 2023), LZ (Aalbers et al., 2023) and the plethora of other direct detection experiments, including those using a similar technique (Adhikari et al., 2021; Coarasa et al., 2024), which is difficult to explain theoretically. There is also a possible hint in favour of dark matter particles from the excess of gamma-rays from the galactic centre found in the FERMI-LAT data (Goodenough & Hooper, 2009; Hooper & Goodenough, 2011), which is also highly debated (see, e.g. Song et al. (2024) for the latest analysis). If confirmed, this would correspond to a 30GeV candidate with very suppressed interactions with nuclei to evade the direct detection limits. A dark matter discovery will take more than two highly debated hints (especially given the plethora of negative searches), so the best way to progress is to exploit the wealth of information coming from astrophysical observations. These observations are particularly useful to probe whether dark matter is collisionless (e.g. Boehm & Schaeffer, 2005; Boehm et al., 2014; Mosbech et al., 2023), and if so, determine the lower bound of the particle mass (e.g. Nadler et al., 2021a). The abundance of negative direct detection searches combined with the flexibility of fits to astronomical data allows for the possibility of a more exotic dark matter particle.
@@
The physical properties of different candidate dark matter particles manifest as macroscopic differences on astronomical scales, which are often expressed phenomenologically as cold and warm models of dark matter. To combat the difficulties in observing dark matter in the Universe, cosmological simulations have been used to quantify which characteristics of a particular dark matter model best fit the observations. Results from early CDM simulations were able to shed light on several dark matter properties, such as the discovery of a universal halo density profile across a wide range of mass scales, and the abundance of intermediate to high mass haloes (Navarro et al., 1996; Springel et al., 2005). However, these early simulations also resulted in several well-known tensions with observations, for example, the missing satellites problem (Klypin et al., 1999; Moore et al., 1999), the core-cusp problem (de Blok, 2010), and the too-big-to-fail problem (Boylan-Kolchin et al., 2012), see Sales et al. (2022) for a recent review. Much research has been dedicated to resolving these tensions by improving the quality of the simulations (e.g. Crain et al., 2015; Kaviraj et al., 2017; Pillepich et al., 2018; Hernández-Aguayo et al., 2022; Schaye et al., 2023). These more advanced simulations in recent years bring higher resolutions with better force accuracy and, importantly, the inclusion and advancement of baryonic models.
@@
The missing satellites problem has been the subject of much examination. State-of-the-art CDM simulations of the 1990s and early 2000s overproduced low-mass dark matter haloes by an order of magnitude compared to expectations based on Milky Way satellite observations of the time. Alternative dark matter models were seen as one possible solution, in particular, warm dark matter (WDM). WDM models introduce a free-steaming scale below which dark matter density fluctuations are erased, thus reducing the number of haloes at high redshift and eventually the number of satellites around Milky Way-like galaxies (Bode et al., 2001; Avila-Reese et al., 2001; Boehm & Schaeffer, 2005). The scale at which these low-mass objects are suppressed is set by the mass of the dark matter particle. A heavier WDM particle would look more similar to CDM, whereas a lighter particle produces a much more significant suppression of low-mass halo formation.
@@
The other accepted solution to the missing satellites problem generally invokes baryonic physics to suppress star formation in low mass haloes, often using the feedback of supernova (Governato et al., 2007; Sawala et al., 2016). The energy generated by a supernova event is transferred back into the gas in the simulations, heating the gas. In low-mass haloes with shallow gravitational potential wells, supernova feedback can cause gas outflows with enough energy to eject gas out of the halo entirely, eventually halting star formation. In larger haloes, supernovae still drive outflows, but the potential wells are generally deep enough to retain the gas, which can be accreted back into the galaxy. More efficient supernova models111Throughout this text, we refer to ‘supernova efficiency’ as the stellar mass fraction converted to supernova, i.e. the parameter in Equation 1 of Springel & Hernquist (2003)., where stars capable of going supernova are formed at higher rates, are able to suppress star formation in more massive galaxies, though as galaxies get larger AGNs become the dominant form of feedback (Croton et al., 2006). Alternatively, reionisation was proposed as a solution to the missing satellites problem (Bullock et al., 2000; Benson et al., 2002), where the photons heat the gas to such a point that low mass haloes can’t cool the gas to form galaxies. Importantly, suppression due to reionisation still allows the formation of low-mass galaxies, but preferentially for haloes that form prior to reionisation; thus, the reionisation simulation parameters can be tuned to match satellite observations.
@@
The tensions between simulations and observations have also been eased in recent years by more complete surveys of the satellite galaxies around the Milky Way (e.g. McConnachie, 2012; McConnachie & Venn, 2020). These surveys have been able to observe more low luminosity galaxies and galaxy remnants (streams). Combining the updated observations with increasingly precise modelling has led authors such as Sales et al. (2022) to conclude that the missing satellites problem can be fully explained within a CDM universe.
@@
However, baryonic solutions to the missing satellites problem come with their own complications. The feedback mechanisms used to reduce low-mass galaxy formation are often conceptually similar but numerically inconsistent between simulations. Supernova feedback alone has an array of implementations across popular codebases, including simply increasing nearby gas temperature (Springel & Hernquist, 2003; Governato et al., 2007), to more complicated kinetic and thermal feedback descriptions (Dalla Vecchia & Schaye, 2008; Chaikin et al., 2023). In reality, there are many forms of feedback, from AGN and supernova to photoheating, which all come with their own implementations within different simulation codes (for a recent review see Vogelsberger et al., 2020). While simulations using a CDM cosmology can solve the missing satellites problem through high efficiency feedback, the solutions are degenerate with each other and not mutually exclusive of WDM. There is still work to be done to find methods to distinguish between dark matter models.
@@
In this paper, we explore a new method for distinguishing the underlying dark matter species in simulations. It is well known that the missing satellites problem can be solved by implementing either a WDM with reduced feedback efficiency or CDM with increased feedback efficiency. Therefore, we combine each dark matter type with different supernova efficiencies to create simulations with similar Milky Way satellite populations. However, supernovae are also critically important in the creation and transportation of dust (Aguirre, 1999; Todini & Ferrara, 2001; Bianchi & Schneider, 2007), and so these different supernova efficiencies are expected to produce different dust properties around the host halo. We explore if these different dust properties can then be used to infer (or at least discriminate) the underlying dark matter model. The techniques we present to analyse the dust provide a proof of concept that can be leveraged in future work with more in-depth analysis of dust in simulations and observations.
@@
This paper is organized as follows. In Section 2, we discuss the codes used and simulation setup. Section 3 explains the simulation results with a discussion following in Section 4. We conclude in Section 5. Throughout this paper we assume a flat Planck18 cosmology (Aghanim et al., 2020), where , , , , =0.840, and .
@@ 2.1 Simulation code setup
For the primary analysis, we perform zoom-in hydrodynamic simulations focusing on Milky Way-like systems. We run cold and warm dark matter, the latter having a mass of 3.5 keV. This is in tension with more stringent observational constraints on the WDM mass (Viel et al., 2013; Nadler et al., 2021a, b; Dekker et al., 2022), but within the limits set by Benito et al. (2020) and Newton et al. (2021). The lighter mass is used to ensure an identifiable difference in the small-scale structure that is wholly attributed to the choice of dark matter.
@@
To generate initial conditions for each cold and WDM simulation, we use music, the Multi-Scale Initial Conditions generator (Hahn & Abel, 2011). We apply the Eisenstein & Hu (1998) fitting function to create the input power spectra for the CDM simulations and assume a truncated power spectrum for the WDM simulations, using the Viel et al. (2005) transfer of the fitting function. The Viel et al. (2005) parameterisation of the transfer function gives us a WDM half mode mass, M⊙, at which point we expect significant divergence between cold and warm dark matter simulations. This work is focused on relative differences, so the fitting functions are accurate enough compared to traditional Boltzmann solvers. music disregards any extra thermal velocity the WDM particles may carry in the initial conditions, which can produce extra spurious haloes if not treated correctly (Klypin et al., 1993; Colín et al., 2008). Using this generalised approach to generate the WDM power spectrum avoids any assumptions about the nature of the particle itself. The initial conditions for all simulations are created with the same random seed.
@@
We run the parent dark matter only and the subsequent zoom-in simulations with gadget-4 (Springel et al., 2021), an -body and SPH code with significant improvements to older versions of gadget. gadget-4 makes use of a TreePM algorithm (Xu, 1995; Bagla, 2002) for the collisionless particles and a pressure-entropy SPH prescription (Springel & Hernquist, 2002) for the gas particles. The gravitational potential is softened with a cubic spline kernel and becomes Newtonian at distances , where epsilon is the simulation force resolution, and the SPH properties are smoothed over 64 nearest neighbours. The baryonic prescriptions implemented with the public release of gadget-4 are discussed in Section 2.3. Finally, we use a friends-of-friends algorithm (Davis et al., 1985) and subfind-hbt (Han et al., 2018; Springel et al., 2021) for halo and subhalo identification.
@@
Both parent and zoom-in simulations are run from redshift to , with a box side length of 65Mpc, noting that the zoom region is 0.1 percent the total volume. The parent simulations have 5123 equal mass dark matter particles, which equates to an approximate mass resolution of and a force resolution of kpc. The zoom-in simulations have a dark matter particle resolution of M⊙ and a gas particle resolution of . More details about the simulations can be found in Table 1. The parent simulations are used to identify the haloes resembling the Milky Way halo to resimulate.
@@
| Box side length | Mpc | Mpc |
| Dark matter mass resolution | M⊙ | M⊙ |
| Baryon mass resolution | NA | M⊙ |
| Stellar mass resolution | NA | M⊙ |
| Force resolution | kpc | pc |
@@ 2.2 Target selection
The halo catalogues produced by subfind-hbt (Springel et al., 2021) are used to identify haloes similar in mass to the Milky Way in the parent cold and WDM simulations. We select haloes with a mass between 1-2 M⊙, where is the mass enclosed with a radius such that the density is 200 times the critical density of the universe, . This is consistent with observed Milky Way masses from Bland-Hawthorn & Gerhard (2016). We then apply an isolation criterion to these haloes. We exclude any potential resimulation target if there is another halo with a mass greater than 0.5 M⊙ within 8Mpc. This ensures that the Milky Way analogue is the largest object in the resimulated region, and that we are not investigating a galaxy undergoing a major merger. We use the same criteria and check both warm and CDM parent simulations to verify the target halo exists in both simulations with similar properties. If there are multiple haloes that fit these criteria, we use the halo closest to the centre of the box.